OpenStudy (anonymous):

Find the quotient . 4/9 / 7/8 ? help plz ?thank u :)

OpenStudy (bahrom7893):

(4/9)*(8/7)

OpenStudy (bahrom7893):

=32/63

OpenStudy (lgbasallote):

go bah and jack!

OpenStudy (anonymous):

4/9 \[\div\] 7/8= 4/9 x 8/7=32/63

OpenStudy (anonymous):